# TAUE: Training-free Noise Transplant and Cultivation Diffusion Model

### v1

@@ Abstract
Despite the remarkable success of text-to-image diffusion models, their output of a single, flattened image remains a critical bottleneck for professional applications requiring layer-wise control. Existing solutions either rely on fine-tuning with large, inaccessible datasets or are training-free yet limited to generating isolated foreground elements, failing to produce a complete and coherent scene. To address this, we introduce the Training-free Noise Transplantation and Cultivation Diffusion Model (TAUE), a novel framework for zero-shot, layer-wise image generation. Our core technique, Noise Transplantation and Cultivation (NTC), extracts intermediate latent representations from both foreground and composite generation processes, transplanting them into the initial noise for subsequent layers. This ensures semantic and structural coherence across foreground, background, and composite layers, enabling consistent, multi-layered outputs without requiring fine-tuning or auxiliary datasets. Extensive experiments show that our training-free method achieves performance comparable to fine-tuned methods, enhancing layer-wise consistency while maintaining high image quality and fidelity. TAUE not only eliminates costly training and dataset requirements but also unlocks novel downstream applications, such as complex compositional editing, paving the way for more accessible and controllable generative workflows.
Figure 1: TAUE introduces a training-free method for layer-wise image generation. By transplanting an intermediate seedling latent from the foreground to the composite generation process, TAUE simultaneously produces a consistent foreground, background, and composite image.
@@ 1 Introduction
Diffusion models have revolutionized creative workflows by enabling the synthesis of photorealistic and intricate images from text prompts (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Rombach et al. 2022; Hu et al. 2024). Yet, this transformative power is constrained by a critical limitation: they typically generate only single-layered, flat images. These flat images severely restrict post-hoc manipulation, as individual elements are inextricably fused. In professional domains such as art, design, and animation, where refining complex compositions is crucial, this lack of layer-wise control presents a significant bottleneck, forcing practitioners to engage in laborious manual segmentation and inpainting.
@@
In response, most existing work has leveraged fine-tuning for layer-wise image generation. These methods attempt to denoise multiple layers simultaneously using masks (Zhang et al. 2023; Huang et al. 2024; Fontanella et al. 2024; Kang et al. 2025) or employ specialized alpha-channel autoencoders (Zhang and Agrawala 2024; Dalva et al. 2024; Huang et al. 2025a; Pu et al. 2025). While effective to some extent, these approaches depend on large-scale, curated datasets and prohibitive training costs. Critically, these datasets are often proprietary or inaccessible, creating a significant barrier to entry that hinders reproducible research and slows community progress. To circumvent this data dependency, training-free approaches have been explored (Quattrini et al. 2024; Morita et al. 2025; Zou et al. 2025). However, these methods focus exclusively on generating isolated foregrounds and do not attempt to produce a corresponding background. This inherently limits them to partial solutions, leaving a critical research gap for a framework capable of layer-wise image generation without requiring fine-tuning.
To address this gap, we propose the novel Training-free Noise Transplant and Cultivation Diffusion Model (TAUE). Inspired by Guo et al. (2024), TAUE introduces a novel mechanism that manipulates this initial noise to achieve semantically coherent, layer-wise image generation. Our core technique, Noise Transplantation and Cultivation (NTC), harvests intermediate latents from the diffusion processes and reuses them to cultivate coherent layers throughout the generation pipeline. First, foreground seedling noise is extracted from the foreground generation and transplanted into the initial noise for the composite generation. Then, background seedling noise is derived from the composite generation and used as the initial noise for background generation. This ensures semantic and structural coherence across layers, enabling consistent, multi-layered outputs—foreground, background, and composite—without requiring fine-tuning or additional datasets. By transplanting and cultivating latents at each stage, TAUE generates a coherent composite scene in which all layers align seamlessly. This approach eliminates the need for expensive fine-tuning and large datasets, while opening up new applications, such as complex compositional editing.
Extensive experiments demonstrate that TAUE achieves state-of-the-art results among training-free methods, improving layer-wise consistency while matching the quality of fine-tuned models. Our contributions are threefold: (1) We introduce TAUE, a novel framework for layered image generation that eliminates the need for fine-tuning and external datasets. (2) We propose NTC, which manipulates ancestral noise of the diffusion process. TAUE achieves natural coherence across layers that previous approaches could only approximate through large-scale training. (3) We demonstrate state-of-the-art performance in training-free layered image generation, enabling a range of previously intractable downstream applications.
Figure 2: The generation process consists of three stages: (1) Foreground Generation, where an object is generated from noise and a seedling latent is extracted; (2) Composite Generation, where the seedling latent is transplanted into a new denoising trajectory to generate a full scene; and (3) Background Generation, where the background is reconstructed separately from the same noise. To ensure spatial and semantic consistency, we introduce an NTC strategy, which constrains the object region with fixed seedling noise during early denoising steps and gradually relaxes this constraint to produce a coherent composite.
@@ 2.1 Initial Noise of Diffusion Model
Diffusion models turn Gaussian noise into structured data through a step-by-step refinement process. Tailoring the initial noise improves image fidelity and alignment, with recent studies focusing on optimal noise selection (Xu, Zhang, and Shi 2024; Wang et al. 2024; Guo et al. 2024; Eyring et al. 2024) and latent optimization (Chen et al. 2024a; Zhou et al. 2024; Qi et al. 2024). Furthermore, identifying optimal noise seeds and focusing on specific noise regions affect image quality and object placement (Ban et al. 2024; Han et al. 2025; Izadi et al. 2025), while noise manipulation enables advanced tasks for layout control (Shirakawa and Uchida 2024; Mao, Wang, and Aizawa 2023b), editing (Mao, Wang, and Aizawa 2023a; Chen et al. 2024b), and video generation (Wu et al. 2024; Morita et al. 2025).
Building on these insights, we propose a novel NTC method that embeds an object’s intermediate latent into the initial latent space. By aligning this noise’s distribution with the Gaussian prior, our method reconstructs the object while generating an entirely new background.
@@
Table 1: Qualitative capability comparison between our TAUE and well-established layer-wise image generation methods such as LayerDiffuse (Zhang and Agrawala 2024), ART (Pu et al. 2025), Alfie (Quattrini et al. 2024). TAUE uniquely enables training-free generation of complete scenes with background synthesis, multi-object disentanglement, semantic harmonization, and layout controllability surpassing both fine-tuned and training-free baselines.
@@ 2.2 Layer-wise Image Generation
Layer-wise image generation is essential for professional applications that require fine-grained compositional control and editing. Fine-tuned approaches typically achieve this using special masks (Zhang et al. 2023; Huang et al. 2024; Fontanella et al. 2024; Kang et al. 2025) or alpha-channel autoencoders (Zhang and Agrawala 2024; Dalva et al. 2024; Huang et al. 2025a; Pu et al. 2025). However, these methods rely heavily on large-scale, proprietary datasets, which limits their practicality and reproducibility.
Training-free alternatives (Morita et al. 2025; Quattrini et al. 2024; Zou et al. 2025) eliminate the need for data collection; however, they fail to produce complete scenes, generating only isolated foregrounds. In contrast, our approach is the first zero-shot, complete solution for layer-wise image generation without fine-tuning or data collection.
@@ 2.3 Layer Decomposition and Composition
An alternative to direct layered generation is post-hoc decomposition, where a flat image is broken into distinct layers. This research focuses on building large-scale datasets to train decomposition models. These datasets may offer either fine-grained part masks, which lack the transparency needed for flexible compositing (Liu et al. 2024), or full RGBA layers (Huang et al. 2025b). However, such datasets are often curated by other models and often fail to capture complex visual effects like soft shadows and reflections, limiting their utility for seamless editing (Burgert et al. 2024; Tudosiu et al. 2024; Yang et al. 2025).
Moreover, beyond decomposition, harmonizing extracted layers into a cohesive image poses a distinct challenge. This step typically requires costly fine-tuning to blend lighting and color—a process that can alter the original appearance of elements (Winter et al. 2024; Zhang, Wen, and Shi 2020; Huang et al. 2025a). These dual challenges—dependence on large datasets and expensive harmonization—underscore fundamental limitations of this pipeline. In contrast, TAUE bypasses these issues by generating coherent foreground, background, and composite layers simultaneously in zero-shot fashion, ensuring consistency from the outset without datasets or post-hoc correction.
@@ 3 TAUE
TAUE utilizes Latent Diffusion Model (LDM) (Rombach et al. 2022) to achieve layer-wise image generation without requiring fine-tuning or additional datasets. In LDM such as Stable Diffusion (Rombach et al. 2022) and SDXL (Podell et al. 2023), the VAE encodes an RGB image into a latent representation $\mathbf{z}\in\mathbb{R}^{4\times H/8\times W/8}$, i.e. four feature channels. We follow the common 0-based indexing convention $(c=0,1,2,3)$ when referring to individual channels.
As shown in Fig. 2 The process of TAUE is divided into three phases: (1) Foreground Generation, (2) Composite Generation with NTC, and (3) Background Generation with NTC. We use three distinct text prompts — $T_{\text{fg}}$, $T_{\text{bg}}$, and $T_{\text{all}}$ — for foreground, background, and composite generation.
  • (1)
    We generate the foreground object on a uniform background to obtain a clean foreground layer $I_{\text{fg}}$3.1). During this step, an intermediate foreground seedling latent $L_{\text{fg}}$ is extracted, encoding structural and semantic features of the object.
  • (2)
    $L_{\text{fg}}$ is transplanted into the initial noise $z_{\text{all},T}$ for composite generation, yielding both the background seedling latent $L_{\text{bg}}$ and the composite scene $I_{\text{all}}$3.2).
  • (3)
@@ 3.1 Foreground Generation
TAUE enables flexible multi-object generation, specifying object positions and sizes with a binary mask. However, uniform binary masks often lead to rigid, box-shaped foregrounds due to their flat probability distribution.
To address this limitation, we introduce a probabilistic mask generation strategy that controls the likelihood of each pixel being included in the mask. Rather than uniformly sampling all pixels within a rectangular region, we define a spatial probability distribution that reflects the spatial importance of pixels within an object’s bounding box, i.e., the region intended for object generation.
Specifically, for a given bounding box centered at $(o_{x},o_{y})$ with width $w$ and height $h$, we define a radially symmetric Gaussian distribution confined to this region:
[…]
Here, $P(x,y)$ serves as an object retention score, indicating how strongly the original noise at pixel $(x,y)$ is preserved in this operation. High values of $P(x,y)$ help maintain the initial noise and thereby promote object synthesis in those areas. The parameter $\sigma$ controls spatial falloff within the bounding box, and $(o_{x},o_{y})$ specifies the center of the object’s intended region. Pixels outside the box are assigned zero probability, ensuring the mask is strictly confined to the object’s designated area.
This retention map is further scaled to a target range $p_{\text{min}},p_{\text{max}}]$ and compared with a randomly sampled value $R(x,y)$ to generate a binary mask:
$M(x,y)=\begin{cases}1&\text{if }R(x,y)>P(x,y)\\ 0&\text{otherwise}\end{cases}.$ (2)
As a result, pixels near the center of the bounding box are less likely to be masked, while those near the edges are more likely to be included in the mask.
Inspired by TKG-DM (Morita et al. 2025), we blend a green background latent vector $C_{\text{gb}}=[0,1,1,0]$ into the initial noise $z_{T}\sim N(0,1)$ within the generated mask $M$:
@@
where $\alpha$ is the mixing parameter controlling the contribution of the green background. Here $C_{\text{gb}}\in\mathbb{R}^{4}$ matches the latent channel count; we set $c=1,2$ (roughly corresponding to the G and B axes after decoding, in an RGB-like layout) to 1 and leave $c=0,3$ at 0. The resulting latent $z_{\text{fg},T}$ is then used in the vanilla denoising process with foreground text prompt $T_{\text{fg}}$ to obtain $I_{\text{fg}}$.
We cache the latent tensor at the crop timestep $t_{\text{crop}}$, which corresponds to a predefined denoising progress ratio $r_{\text{crop}}$. Specifically, $t_{\text{crop}}$ is defined as the timestep at which $r_{\text{crop}}$ of the total denoising process has been completed, measured along the reverse time axis of the scheduler. We then define:
@@
where $t_{\text{crop}}=\lfloor T\cdot(1-r_{\text{crop}})\rfloor$. This cached latent, hereafter called the seedling latent, encodes foreground geometry and semantics and will be re-injected in the subsequent stages.
3.2 Composite Generation with Noise Transplantation and Cultivation (NTC)
Just as a seedling can shape the ecosystem it grows in, transplanting latent noise from foreground into new generative processes can guide the formation of coherent visual scenes. In this phase, we generate a coherent composite scene by transplanting the seedling latent $L_{\text{fg}}$ from the foreground phase and cultivating it within the diffusion process. This is guided by an object mask and reinforced through spatially-aware attention and noise control mechanisms.
Object Region Mask   We begin by localizing the object region in the latent space using two complementary signals: latent channel activations and cross-attention maps. During foreground generation, a green latent vector $C_{\text{gb}}$ is injected into the background region of the initial latent, elevating activations in channels $c=1$ and $c=2$. The object region, unaffected by this injection, exhibits low activation in these channels. We construct a smooth activation map as:

### v2

@@ Abstract
Despite the remarkable success of text-to-image diffusion models, their output of a single, flattened image remains a critical bottleneck for professional applications requiring layer-wise control. Existing solutions either rely on fine-tuning with large, inaccessible datasets or are training-free yet limited to generating isolated foreground elements, failing to produce a complete and coherent scene. To address this, we introduce the Training-free Noise Transplantation and Cultivation Diffusion Model (TAUE), a novel framework for layer-wise image generation that requires neither fine-tuning nor additional data. TAUE embeds global structural information from intermediate denoising latents into the initial noise to preserve spatial coherence, and integrates semantic cues through cross-layer attention sharing to maintain contextual and visual consistency across layers. Extensive experiments demonstrate that TAUE achieves state-of-the-art performance among training-free methods, delivering image quality comparable to fine-tuned models while improving inter-layer consistency. Moreover, it enables new applications, such as layout-aware editing, multi-object composition, and background replacement, indicating potential for interactive, layer-separated generation systems in real-world creative workflows.
Figure 1: TAUE introduces a training-free method for layer-wise image generation, enabling independent foreground–background synthesis with a coherent composite output. Our method further enables applications such as layout and size control, disentangled multi-object generation, and background replacement, broadening the creative and practical capabilities of real-world workflows.
@@ 1 Introduction
Diffusion models have revolutionized creative workflows by enabling the synthesis of photorealistic and intricate images from text [16, 36, 34, 17]. Yet, this transformative power is constrained by a critical limitation: they typically generate only single-layered, flat images. These flat images severely restrict post-hoc manipulation, as individual elements are inextricably fused. In professional domains such as art, design, and animation, where refining complex compositions is crucial, this lack of layer-wise control presents a significant bottleneck, forcing practitioners to engage in laborious manual segmentation and inpainting.
@@
In response, most existing work has leveraged fine-tuning for layer-wise image generation. These methods attempt to denoise multiple layers simultaneously using masks [47, 20, 12, 22] or employ specialized alpha-channel autoencoders [44, 9, 18, 31]. While effective to some extent, these approaches depend on large-scale, curated datasets and prohibitive training costs. Critically, these datasets are often proprietary or inaccessible, creating a significant barrier to entry that hinders reproducible research and slows community progress. To circumvent this data dependency, training-free approaches have been explored [33, 28, 49]. However, these methods focus exclusively on generating isolated foregrounds and do not attempt to produce a corresponding background. This inherently limits them to partial solutions, leaving a critical research gap for a framework capable of layer-wise image generation without fine-tuning.
To address this gap, we propose the novel Training-free Noise Transplant and Cultivation Diffusion Model (TAUE) [^1] Taue is the Japanese term for transplanting rice seedlings, symbolizing the nurturing and cultivation process, which parallels our model’s core concept of cultivating latent representations., independently generates foreground and background, and coherent composite images. Instead of regenerating all layers from scratch, TAUE reuses intermediate latents extracted during the denoising process and transplants them into the initial noise of subsequent generations, enabling structural continuity across layers. Moreover, TAUE introduces cross-layer attention sharing, enabling the exchange of semantic information between the foreground and background to maintain scene-level coherence and context alignment. As a result, TAUE produces consistent multi-layered outputs and supports complex compositional editing and interactive layer-separated generation without any fine-tuning or additional data.
Extensive experiments demonstrate that TAUE achieves state-of-the-art results among training-free methods, improving layer-wise consistency while matching the quality of fine-tuned models. Our contributions are as follows:
  • We introduce TAUE, a training-free framework for layer-wise image generation that produces coherent foreground, background, and composite images without any fine-tuning or external datasets.
  • TAUE combines latent transplantation, which embeds structural information from intermediate latent into the initial noise, with cross-layer attention sharing to propagate semantic cues, achieving consistent and contextually aligned multi-layer generation.
  • Our method outperforms existing models, delivering results that match or surpass fine-tuned alternatives while maintaining lower computational costs.
  • TAUE enables low-cost, practical, and interactive applications such as complex compositional editing and layer-separated image generation.
Figure 2: TAUE consists of three stages: (1) Foreground Generation, where an object is generated from noise and an intermediate latent is extracted; (2) Composite Generation, where the extracted latent is transplanted into a new denoising trajectory to synthesize a coherent foreground–background scene through cross-layer attention sharing; and (3) Background Generation, which mirrors the composite phase with the object mask inverted to refine the background region. This three-stage design enables TAUE to maintain spatial and semantic coherence across layers, producing unified, multi-layered image generation without any fine-tuning or additional data.
@@ 2.1 Initial Noise of Diffusion Model
Diffusion models turn Gaussian noise into structured data through a step-by-step refinement process. Tailoring the initial noise improves image fidelity and alignment, with recent studies focusing on optimal noise selection [41, 38, 13, 11, 2] and latent optimization [6, 48, 32, 10]. Furthermore, identifying optimal noise seeds and focusing on specific noise regions affect image quality and object placement [3, 14, 21], while noise manipulation enables advanced tasks for layout control [35, 27], editing [26, 7], and video generation [40, 28].
Building on these insights, we propose embedding the intermediate latent into the initial noise of the diffusion process. This allows the model to preserve the structural information of the object such as its shape and spatial configuration while reconstructing the image with improved layer-wise consistency and generating entirely new content guided by the text condition.
Table 1: Qualitative capability comparison between our TAUE and well-established layer-wise image generation methods such as LayerDiffuse [44], ART [31], Alfie [33]. TAUE uniquely enables training-free generation of complete scenes with background synthesis, multi-object disentanglement, semantic harmonization, and layout controllability surpassing both fine-tuned and training-free baselines.
@@ 2.2 Layer-wise Image Generation
Layer-wise image generation is essential for professional applications that require fine-grained compositional control and editing. [8, 29, 4] Fine-tuned approaches typically achieve this using special masks [47, 20, 12, 22] or alpha-channel autoencoders [44, 9, 18, 31]. However, these methods rely heavily on large-scale, proprietary datasets, which limits their practicality and reproducibility.
Training-free alternatives [28, 33, 49] eliminate needs for data collection; however, they fail to produce complete scenes, generating only isolated foregrounds. In contrast, our approach is the first zero-shot, complete layer-wise image generation without fine-tuning or external data.
@@ 2.3 Layer Decomposition and Composition
An alternative to direct layered generation is post-hoc decomposition, where a flat image is broken into distinct layers. This research focuses on building large-scale datasets to train decomposition models. These datasets may offer either fine-grained part masks, which lack the transparency needed for flexible compositing [25], or full RGBA layers [19]. However, such datasets are often curated by other models and often fail to capture complex visual effects like soft shadows and reflections, limiting their utility for seamless editing [4, 37, 42].
Moreover, beyond decomposition, harmonizing extracted layers into a cohesive image poses a distinct challenge. This step typically requires costly fine-tuning to blend lighting and color, a process that can alter the original appearance of elements [39, 43, 18]. These dual challenges, dependence on large datasets and expensive harmonization, underscore fundamental limitations of this pipeline. In contrast, TAUE bypasses these issues by generating coherent foreground, background, and composite layers simultaneously in zero-shot fashion, ensuring consistency from the outset without datasets or post-hoc correction.
@@ 3 TAUE
3.1 Overall
TAUE utilizes Latent Diffusion Model (LDM) [34] to achieve layer-wise image generation without requiring fine-tuning or additional datasets. As shown in Fig. 2, TAUE is divided into three phases: (1) Foreground Generation, (2) Composite Generation, and (3) Background Generation. We use three distinct text prompts — $T_{\text{fg}}$, $T_{\text{bg}}$, and $T_{\text{all}}$ — for foreground, background, and composite generation.
  1. (1)
    We generate the foreground object on a uniform background to obtain a clean foreground layer $I_{\text{fg}}$3.2). During this step, an intermediate foreground latent $L_{\text{fg}}$ is extracted, encoding structural features of the object.
  2. (2)
    $L_{\text{fg}}$ is transplanted into the initial noise $z_{\text{all},T}$ for composite generation, yielding both the intermediate background latent $L_{\text{bg}}$ and the composite scene $I_{\text{all}}$3.3).
  3. (3)
@@ 3.2 Foreground Generation
TAUE first generates an isolated foreground object according to the foreground prompt $T_{\text{fg}}$. Inspired by TKG-DM [28], which demonstrated that adding channel-wise biases to the initial Gaussian noise $z_{T}\in\mathbb{R}^{4\times H/8\times W/8}$ enables color-controlled image synthesis, we apply a similar strategy to guide the background color in the latent space. Specifically, we blend a green background latent vector $C_{\text{gb}}=[0,1,1,0]$ ($C_{\text{gb}}\in\mathbb{R}^{4\times H/8\times W/8}$) into the initial noise $z_{T}$ within a spatial mask $M$:
@@
where $\alpha$ controls the blending strength of the background color. The mask $M$ defines the spatial region for foreground synthesis and is detailed in the following subsection (§3.2.1). The resulting latent $z_{\text{fg},T}$ is then denoised using the foreground prompt $T_{\text{fg}}$ to produce the isolated foreground image $I_{\text{fg}}$.
3.2.1 Layout Specification
Existing methods typically employ a Gaussian or binary box-shaped mask to localize the foreground region as a spatial mask $M$. However, this approach often produces artifacts along the mask boundary, as the generated content tends to conform too closely to the mask shapes. To overcome this limitation, we redefine $M$ as a probabilistic layout mask that decouples object generation from mask edges by introducing a spatially weighted sampling scheme.
Given a bounding box centered at $(o_{x},o_{y})$ with width $w$ and height $h$, we define a radially symmetric Gaussian distribution confined to this region:
$P(x,y)=\begin{cases}\exp\!\left(-\frac{1}{2\sigma^{2}}\left[\left(\frac{x-o_{x}}{w/2}\right)^{2}+\left(\frac{y-o_{y}}{h/2}\right)^{2}\right]\right),\\ \quad\text{if }|x-o_{x}|\leq w/2,\ |y-o_{y}|\leq h/2,\\ 0,\quad\text{otherwise}.\end{cases}$ (2)
where $P(x,y)$ serves as an object retention score, indicating how strongly the original noise at each pixel is preserved. Larger values promote stable synthesis near the center, while lower values allow flexible adaptation near edges. After rescaling $P(x,y)$ to $p_{\min},p_{\max}]$, we generate a binary mask by stochastic comparison with a random matrix $R(x,y)$:
$M(x,y)=\begin{cases}1&\text{if }R(x,y)>P(x,y),\\ 0&\text{otherwise.}\end{cases}$ (3)
This probabilistic masking allows smooth transitions at object boundaries, solving artifacts tied to the mask contour and enabling flexible control over position and scale. The resulting mask $M$ is then used for latent blending during foreground generation (§3.2), ensuring that the generated foreground follows the intended layout while remaining structurally coherent.
3.2.2 Intermediate Latent Extraction
During denoising, we cache the intermediate latent tensor at a specific timestep $t_{\text{crop}}$, corresponding to a predefined denoising ratio $r_{\text{crop}}$. Formally,
@@ 3.2.2 Intermediate Latent Extraction
where $t_{\text{crop}}=\lfloor T\cdot(1-r_{\text{crop}})\rfloor$. We refer to this cached latent, which encodes the geometric and semantic structure of the foreground object and will be transplanted in subsequent stages to guide composite and background generations.
Figure 4: Qualitative comparison of layer-wise image generation. For each case, we show the foreground, background, and composite image generated by LayerDiffuse [44], Alfie [33], and our method. TAUE consistently produces spatially aligned and semantically coherent multi-layer outputs, achieving realistic integration of foreground and background without fine-tuning or inpainting.
3.3 Composite Generation
Just as a seedling can shape the ecosystem it grows in, transplanting latent noise from foreground into new generative processes can guide the formation of coherent visual scenes. In this phase, we generate a coherent composite scene by transplanting the intermediate latent $L_{\text{fg}}$ from the foreground phase and cultivating it within the diffusion process. This is guided by an object mask and reinforced through spatially-aware attention and noise control mechanisms.
3.3.1 Object Region Mask
We begin by localizing the object region in the latent space using two complementary signals: latent channel activations and cross-attention maps. During foreground generation, a green latent vector $C_{\text{gb}}$ is injected into the background region of the initial latent, elevating activations in channels $c=1$ and $c=2$. The object region, unaffected by this injection, exhibits low activation in these channels. We construct a smooth activation map as: